Grade 12Mechanics

a bullet of mass 10g travelling at a speed of 500 m/s strikes a block of mass 2 kg, which is suspended by a string of length 5 m. The centre of gravity of the block is found to rise a vertical distance of 0.2 m . what is speed of bullet when it emerges from block.

lavish
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
At first use the principle of conservation of mechanical energy...i.e..mgh=1/2mv*v....put the values...the value of v will be used in the next step...conservation of linear momentum....(m1*u1)+(m2*u2)=(m1*v1)+(m2*v2)......it wiil give the value of v1...which will be ur answer....
 
Regards
Arun